# A Prompt Learning Framework for Source Code Summarization

### v1

@@ 4.1. Overview
Figure 2. Overview of PromptCS
4.2. Code Embedding Generation
@@
4.3. Prompt Embedding Generation
The function of the prompt embedding is to instruct the LLM to generate code summaries. As shown in Figure 2(1), the prompt embedding $\bm{e}^{P}$ is generated through the prompt encoder. The input of the prompt encoder is a pseudo prompt consisting of $n$ learnable tokens, denoted $p=\{t_{0},t_{1},\dots,t_{n}\}$, where $t_{i}$ refers to the $i$-th pseudo token. The role of pseudo tokens is to serve only as placeholders without carrying any actual meaning, and to indicate the length of the prompt used by the prompt encoder to generate the prompt embedding. The prompt encoder is a DL model capable of mapping $p$ to a continuous sequence of numbers, which is then input into the embedding layer to generate $\bm{T}=\{\bm{t}_{0},\bm{t}_{1},\dots,\bm{t}_{n}\}$, where $\bm{t}_{i}$ is a trainable embedding tensor. This allows us to discover more effective continuous prompts that go beyond the original vocabulary of the LLM. Then, the prompt encoder applies a Bidirectional Long-short Term Memory network (BiLSTM), with a ReLU activated two-layer multilayer perceptron (MLP) to produce the real prompt embeddings $\bm{e}^{P}=\{\bm{e}_{0},\bm{e}_{1},\dots,\bm{e}_{n}\}$. The value $\bm{e}_{i}$ at the $i$-th position of $\bm{e}^{P}$ is calculated as follows:
(1) $\footnotesize\bm{h}_{i}^{f}=LSTM(\bm{h}_{i-1}^{f}),\;\bm{h}_{i}^{b}=LSTM(\bm{h}_{i+1}^{b}),\;\bm{e}_{i}=MLP([\bm{h}_{i}^{f},\bm{h}_{i}^{b}])$
@@ 4.3. Prompt Embedding Generation
4.4. Fusion Embedding Generation
Similar to how humans concatenate the discrete prompt with the code snippet, PromptCS can also concatenate the continuous prompt generated by the prompt agent (i.e., prompt embedding) with the embedding of the code snippet in various ways. In order not to destroy the integrity of the code itself, a common human practice is to concatenate discrete prompts in front or back of the code snippet (Tian et al., 2023; Sun et al., 2023c). The left side of Figure 3 shows an example of concatenating a discrete prompt and a code snippet (denoted as ¡code¿), where (a) and (b) represent two modes respectively: front-end mode, concatenating the discrete prompt in front of the code snippet; and back-end mode, concatenating the discrete prompt in front of the code snippet. In this paper, in addition to following the aforementioned two modes, we also try a new mode called the two-end mode. In the two-end mode, the prompt embedding is split into two parts that will be concatenated in the front and back of the code embedding, respectively. This split is easy to implement. For example, we can split $\bm{e}^{P}=\{\bm{e}_{0},\bm{e}_{1},\dots,\bm{e}_{n}\}$ into $\bm{e}^{P}_{1}=\{\bm{e}_{0},\bm{e}_{1},\dots,\bm{e}_{i}\}$ and $\bm{e}^{P}_{2}=\{\bm{e}_{i+1},\dots,\bm{e}_{n}\}$. The right side of Figure 3 shows an example of concatenating a prompt embedding and a code embedding where (e), (f), and (g) showcase three concatenation modes, respectively.
Figure 3. Example of prompt and code concatenation
4.5. Model Training
@@
(2) $\mathcal{L}(\Theta)=-\sum_{i=1}^{C}y_{i}log\frac{exp(\hat{y_{i}})}{\sum_{j=1}^{C}exp(\hat{y_{j}})}$
@@ 5. Evaluation and Analysis
5.1. Experimental Setup
5.1.1. Dataset
We evaluate PromptCS on the CodeSearchNet (CSN) corpus (Husain et al., 2019), which is an extensive collection of code snippets accompanied by their comments written in six programming languages (including Go, Java, JavaScript, PHP, Python, and Ruby). This corpus is most commonly used in studying code summarization (Guo et al., 2022; Zhou et al., 2022; Ahmed and Devanbu, 2022b; Sun et al., 2023a, c). Considering that the original CSN corpus contains some low-quality data (Lu et al., 2021), we use the dataset from the CodeXGLUE (Lu et al., 2021) code-to-text docstring generation task, which is built upon the CSN corpus and excludes defective data samples.
5.1.2. Evaluation Metrics
We leverage four metrics in the evaluation, including BLEU (Papineni et al., 2002), METEOR (Banerjee and Lavie, 2005), ROUGE (Lin, 2004), and SentenceBERT (Reimers and Gurevych, 2019). These metrics are widely used in code summarization (Iyer et al., 2016; Wan et al., 2018; Hu et al., 2018b; Zhang et al., 2020; Wang et al., 2022; Wu et al., 2021; Sun et al., 2023a).
BLEU, the abbreviation for BiLingual Evaluation Understudy (Papineni et al., 2002), is a variant of precision metric, which calculates the similarity by computing the n-gram precision of a generated summary to the ground-truth summary, with a penalty for the overly short length. In this paper, we follow (Wu et al., 2021; Sun et al., 2023a) and show the standard BLEU score which provides a cumulative score of 1-, 2-, 3-, and 4-grams.
METEOR (Banerjee and Lavie, 2005) is introduced to address the concerns of using BLEU. METEOR combines n-gram precision and n-gram recall by taking their harmonic mean to compute a measure of similarity.
ROUGE-L, is a variant of ROUGE (Recall-oriented Understudy for Gisting Evaluation) (Lin, 2004). ROUGE-L is computed based on the longest common subsequence (LCS).
SentenceBERT (Haque et al., 2022). Unlike the above three metrics that mainly calculate the textual similarity between the ground truth summaries and the generated summaries, SentenceBERT measures the semantic similarity. SentenceBERT first converts the two compared summaries into embeddings in a unified vector space, and then uses cosine similarity to represent the semantic similarity between them.
@@ 5.1.3. Base LLMs
We conduct experiments on four popular autoregressive LLMs, including three open-source LLMs: PolyCoder (Xu et al., 2022), CodeGen-Multi (Nijkamp et al., 2023), StarCoderBase (Li et al., 2023), and a commercial LLM: ChatGPT (OpenAI, 2022).
@@
PolyCoder (Xu et al., 2022) is released by researchers from Carnegie Mellon University. It is trained using the GPT NeoX toolkit on 249GB of code from 12 programming languages, and is available in three sizes: 160M, 0.4B, and 2.7B.
@@
CodeGen-Multi (Nijkamp et al., 2023) is released by Salesforce, which is a member of an autoregressive language models family for program synthesis. In our experiments, we use three sizes of CodeGen-Multi: 350M, 2B, and 6B.
@@
StarCoderBase (Li et al., 2023) is jointly released by Hugging Face and ServiceNow in 2023. It is trained on 80+ programming languages from The Stack (Kocetkov et al., 2022). And the renowned StarCoder is a fine-tuned version of StarCoderBase, trained on an additional 35 billion Python tokens. Considering that StarCoder is only available in a 16B version, which may not be suitable for conducting comprehensive comparison experiments. In our experiments, we use three sizes of StarCoderBase: 1B, 3B, and 7B.
@@
ChatGPT (OpenAI, 2022) is possibly the most powerful LLM. To conduct experiments with its assistance, we utilize the OpenAI API, which is powered by a diverse set of models with different capabilities, such as GPT-3.5, DALL·E and Whisper. In our experiments, we use the gpt-3.5-turbo model, which distinguishes itself as the most capable and cost-effective option within the GPT-3.5 family.
5.1.4. Experimental Settings
@@
At the training stage, we set the mini-batch size to 16, and the learning rate to 5e-5. We employ the AdamW optimizer (Loshchilov and Hutter, 2019) along with a linear learning rate scheduler, following the default setup recommended by Hugging Face. To ensure comprehensive training, we employ an early stopping strategy based on the best validation BLEU to control the number of training epochs for the model. The early stopping patience is set to 4. The input sequences are padded to the maximum length with special tokens, which are selected based on the vocabulary of the LLM. To fine-tune StarCoderBase-7B and CodeGen-Multi-6B on a single A800, we utilize DeepSpeed 0.12.2. We set the Zero Redundancy Optimizer (ZeRO) to ZeRO-3 and enable the offloading of optimizer computation to CPU. All models are implemented using the PyTorch 2.1.0 framework with Python 3.8. All experiments are conducted on a server equipped with one Nvidia A800 GPU with 80 GB memory, running on Ubuntu 20.04.4.
@@
Table 2. Effectiveness of PromptCS on the CSN-Java dataset. Zero-shot: Instruction Prompting with zero-shot learning; Few-shot: Instruction Prompting with few-shot learning; $\mathcal{B}$: BLEU; $\mathcal{M}$: METEOR; $\mathcal{R}$: ROUGE-L; $\mathcal{S}$: SentenceBERT.
@@
7B 9.47 8.05 20.10 30.24 16.65 12.91 33.67 58.31 $*{}^{*}$${}^{*}$${}^{*}$${}^{*}$ 21.86 15.08 41.25 63.21
ChatGPT 9.57 14.89 20.75 54.11 14.68 15.97 34.61 59.89 ${}^{\dagger}$$^{\dagger}$$^{\dagger}$$^{\dagger}$$^{\dagger}$$^{\dagger}$$^{\dagger}$$^{\dagger}$
  • $*{}^{*}$ StarCoderBase-7B and CodeGen-Multi-6B are too large for effective fine-tuning.
@@
  • ${}^{\dagger}$ ChatGPT is a non-open source model and cannot be fine-tuned. We also cannot apply PromptCS to it.
5.2. Experimental Results
5.2.1. RQ1: Effectiveness of PromptCS
Instruct Prompting (zero-shot). This scheme directly uses human-written instructions to prompt LLMs to generate summaries for given code snippets. For StarCoderBase, CodeGen-Multi, and PolyCoder, we utilize the human-written prompt provided in the work (Bhattacharya et al., 2023). For ChatGPT, we utilize the human-written prompt provided in the work (Sun et al., 2023c). Table 3 presents the two human-written prompts mentioned above.
Table 3. Manual prompts used to guide LLMs for code summarization
LLM Prompt Instruction
StarCoderBase //Human: You are a helpful code summarizer. Please describe in simple english the purpose of the following Java code snippet: ¡code¿ //Assistant:
CodeGen-Multi
PolyCoder
ChatGPT Please generate a short comment in one sentence for the following function: ¡code¿
Instruct Prompting (few-shot). In this scheme, in addition to providing the hand-written prompts used in the zero-shot setting, we also provide a few examples that demonstrate the nature of the task to let LLMs perform few-shot learning. We follow (Ahmed and Devanbu, 2022b; Bhattacharya et al., 2023) and provide 10 examples in the few-shot setting. Each example is a pair of $\langle code\;snippet,summary\rangle$ randomly selected from the training set. In practice, we directly leverage the 10 examples provided by Ahmed et al. in their GitHub repository (Ahmed and Devanbu, 2022a), since we use the same experimental dataset (i.e., the CSN corpus).
@@ 5.2.1. RQ1: Effectiveness of PromptCS
2) Results: Table 2 shows the performance of our PromptCS and baselines in terms of the four evaluation metrics, i.e., BLEU, METEOR, ROUGE-L, and SentenceBERT. Observe that task-oriented fine-tuning performs best on the LLM PolyCoder in terms of scores of the four metrics, followed by PromptCS, instruction prompting (few-short), and instruction prompting (zero-shot). On the LLMs CodeGen-Multi and StarCoderBase, PromptCS is overall comparable to task-oriented fine-tuning, and both are significantly better than the two instruction prompting schemes. PromptCS even outperforms task-oriented fine-tuning on some LLMs, such as CodeGen-Multi-2B (row 7), StarCoderBase-1B (row 9) and -3B (row 10). We also conduct instruction prompting schemes on a commercial LLM ChatGPT, and its performance is shown in the last row of Table 2. Observe that the performance of both instruction prompting schemes on ChatGPT is better than that on the other three LLMs in all four metrics, but is also not satisfactory compared to task-oriented fine-tuning and PromptCS. Although this comparison is unreasonable, we have reason to believe that if ChatGPT is open source, the task-oriented fine-tuning scheme and PromptCS will significantly improve its code summarization capabilities compared to the instruction prompting schemes.
@@
Table 4. Training time of Fine-tuning and PromptCS
@@
CodeGen-Multi 350M 20h:06m:44s 15h:01m:39s
2B 81h:08m:24s 58h:43m:29s
6B $*{}^{*}$ 88h:03m:40s
PolyCoder 160M 12h:25m:19s 09h:41m:45s
0.4b 22h:18m:34s 14h:32m:49s
2.7b 81h:54m:24s 27h:41m:15s
StarCoderBase 1B 50h:06m:52s 29h:17m:32s
3B 82h:26m:21s 43h:51m:44s
7B $*{}^{*}$ 67h:21m:31s
  • $*{}^{*}$tarCoderBase-7B and CodeGen-Multi-6B are too large for effective fine-tuning.
    Here is the single-epoch training time for it.
In addition, we also compare the training costs between the task-oriented scheme and PromptCS in terms of training time. Table 4 lists the time costs incurred during training on three LLMs. Observe that compared to the task-oriented scheme, PromptCS noticeably requires less time on all LLMs, which can be attributed to the fact that PromptCS is non-invasive to LLMs and does not update the parameters of the LLMs during training. It should be noted that when the model size of the LLM is very large (e.g., StarCoderBase-7B), the task-oriented fine-tuning will be very time-consuming and become unacceptable if the available training resources (e.g., the performance and number of GPU devices) are few.
@@
Table 5. Effects of prompt length and concatenation mode on PromptCS. $\mathcal{B}$: BLEU; $\mathcal{M}$: METEOR; $\mathcal{R}$: ROUGE-L; $\mathcal{S}$: SentenceBERT.
@@
5.2.2. RQ2. Influence of key configurations on PromptCS
@@
Table 6. Performance of PromptCS when building the prompt encoder on different network architectures. We bold the metric scores where the prompt encoder built on BiLSTM/Transformer performs better.
@@
5.2.3. RQ3. Influence of the network architecture used in the prompt encoder on PromptCS
@@
5.2.4. RQ4. Influence of the training data size on PromptCS
In this paper, we also analyze the impact of training data size on the effectiveness of PromptCS. To disclose this impact, we commence with a small-scale training set and systematically augment the number of training samples. The smaller training sets are randomly sampled from the complete training data. The considered training set sizes encompass 100, 1000, 10000, and 164923 (complete training data). In this experiment, we also uniformly utilize StarCoderBase-1B as the base LLM and set the prompt length and the concatenation mode to 100 and the back-end mode, respectively.
Figure 4. Training costs for different prompt lengths
Figure 4. Training costs for different prompt lengths
Figure 5. Influence of training set size on PromptCS
Figure 5 showcases the experimental results, where the x-axis denotes the training set size, and the y-axes denote BLEU, METEOR, and ROUGE-L on the left and SentenceBERT on the right. Observe that with an increase in the size of the training set, the scores of PromptCS on the four metrics will increase, but the increase is not obvious. And PromptCS trained with 100 samples performs comparably to PromptCS trained with 164,923 samples. This underscores PromptCS’s superior adaptability and generalization capabilities on a small-scale dataset. It holds profound practical significance for reducing training costs, improving training efficiency, and obtaining satisfactory performance in environments with limited data. It holds profound practical significance for reducing training costs, improving training efficiency, and obtaining satisfactory performance in environments with limited data.
@@
Table 7. Effectiveness of PromptCS in different programming languages. $\mathcal{B}$: BLEU; $\mathcal{M}$: METEOR; $\mathcal{R}$:ROUGE-L; $\mathcal{S}$: SentenceBERT. We bold the metric scores where Task-oriented Fine-turning/PromptCS performs better.
@@
Task-oriented Fine-turning PromptCS Task-oriented Fine-turning PromptCS
@@
5.2.5. RQ5. Effectiveness in other programming languages

### v2

@@ 3.1 Overview
Figure 2: Overview of PromptCS
3.2 Code Embedding Generation
@@
3.3 Prompt Embedding Generation
The function of the prompt embedding is to instruct the LLM to generate code summaries. As shown in Figure 2(1), the prompt embedding $\bm{e}^{P}$ is generated through the prompt encoder. The input of the prompt encoder is a pseudo prompt consisting of $n$ learnable tokens, denoted $p=\{t_{0},t_{1},\dots,t_{n}\}$, where $t_{i}$ refers to the $i$-th pseudo token. The role of pseudo tokens is to serve only as placeholders without carrying any actual meaning, and to indicate the length of the prompt used by the prompt encoder to generate the prompt embedding. The prompt encoder is a DL model capable of mapping $p$ to a continuous sequence of numbers, which is then input into the embedding layer to generate $\bm{T}=\{\bm{t}_{0},\bm{t}_{1},\dots,\bm{t}_{i},\dots,\bm{t}_{n}\}$, where $\bm{t}_{i}$ is a trainable embedding tensor. This allows us to discover more effective continuous prompts that go beyond the original vocabulary of the LLM. Then, the prompt encoder applies a Bidirectional Long-short Term Memory network (BiLSTM), with a ReLU activated two-layer multilayer perceptron (MLP) to produce the real prompt embeddings $\bm{e}^{P}=\{\bm{e}_{0},\bm{e}_{1},\dots,\bm{e}_{n}\}$. The value $\bm{e}_{i}$ at the $i$-th position of $\bm{e}^{P}$ is calculated as follows:
$\footnotesize\bm{h}_{i}^{f}=LSTM(\bm{h}_{i-1}^{f}),\;\bm{h}_{i}^{b}=LSTM(\bm{h}_{i+1}^{b}),\;\bm{e}_{i}=MLP([\bm{h}_{i}^{f},\bm{h}_{i}^{b}])$ (1)
@@ 3.3 Prompt Embedding Generation
3.4 Fusion Embedding Generation
Similar to how humans concatenate the discrete prompt with the code snippet, PromptCS can also concatenate the continuous prompt generated by the prompt agent (i.e., prompt embedding) with the embedding of the code snippet in various ways. In order not to destroy the integrity of the code itself, a common human practice is to concatenate discrete prompts in front or back of the code snippet [15, 5]. The left side of Figure 3 shows an example of concatenating a discrete prompt and a code snippet (denoted as ¡code¿), where (a) and (b) represent two modes respectively: front-end mode, concatenating the discrete prompt in front of the code snippet; and back-end mode, concatenating the discrete prompt in front of the code snippet. In this paper, in addition to following the aforementioned two modes, we also try a new mode called the two-end mode. In the two-end mode, the prompt embedding is split into two parts that will be concatenated in the front and back of the code embedding, respectively. This split is easy to implement. For example, we can split $\bm{e}^{P}=\{\bm{e}_{0},\bm{e}_{1},\dots,\bm{e}_{n}\}$ into $\bm{e}^{P}_{1}=\{\bm{e}_{0},\bm{e}_{1},\dots,\bm{e}_{i}\}$ and $\bm{e}^{P}_{2}=\{\bm{e}_{i+1},\dots,\bm{e}_{n}\}$. The right side of Figure 3 shows an example of concatenating a prompt embedding and a code embedding where (e), (f), and (g) showcase three concatenation modes, respectively.
Figure 3: Example of prompt and code concatenation
3.5 Model Training
@@
$\mathcal{L}(\Theta)=-\sum_{i=1}^{C}y_{i}log\frac{exp(\hat{y_{i}})}{\sum_{j=1}^{C}exp(\hat{y_{j}})}$ (2)
@@ 4 Evaluation and Analysis
4.1 Experimental Setup
4.1.1 Dataset
We evaluate PromptCS on the CodeSearchNet (CSN) corpus [60], which is an extensive collection of code snippets accompanied by their comments written in six programming languages (including Go, Java, JavaScript, PHP, Python, and Ruby). This corpus is most commonly used in studying code summarization [57, 61, 28, 3, 5]. Considering that the original CSN corpus contains some low-quality data [62], we use the dataset from the CodeXGLUE [62] code-to-text docstring generation task, which is built upon the CSN corpus and excludes defective data samples.
4.1.2 Evaluation Metrics
We leverage four metrics in the evaluation, including BLEU [63], METEOR [64], ROUGE [65], and SentenceBERT [66]. These metrics are widely used in code summarization [67, 68, 6, 11, 69, 70, 3].
BLEU, the abbreviation for BiLingual Evaluation Understudy [63], is a variant of precision metric, which calculates the similarity by computing the n-gram precision of a generated summary to the ground-truth summary, with a penalty for the overly short length. In this paper, we follow [70, 3] and show the standard BLEU score which provides a cumulative score of 1-, 2-, 3-, and 4-grams.
METEOR [64] is introduced to address the concerns of using BLEU. METEOR combines n-gram precision and n-gram recall by taking their harmonic mean to compute a measure of similarity.
ROUGE-L, is a variant of ROUGE (Recall-oriented Understudy for Gisting Evaluation) [65]. ROUGE-L is computed based on the longest common subsequence (LCS).
SentenceBERT [71]. Unlike the above three metrics that mainly calculate the textual similarity between the ground truth summaries and the generated summaries, SentenceBERT measures the semantic similarity. SentenceBERT first converts the two compared summaries into embeddings in a unified vector space, and then uses cosine similarity to represent the semantic similarity between them.
@@ 4.1.3 Base LLMs
We conduct experiments on four popular autoregressive LLMs, including three open-source LLMs: PolyCoder [26], CodeGen-Multi [25], StarCoderBase [24], and a commercial LLM: ChatGPT [21].
@@
PolyCoder [26] is released by researchers from Carnegie Mellon University. It is trained using the GPT NeoX toolkit on 249GB of code from 12 programming languages, and is available in three sizes: 160M, 0.4B, and 2.7B.
@@
CodeGen-Multi [25] is released by Salesforce, which is a member of an autoregressive language models family for program synthesis. We use three sizes of CodeGen-Multi: 350M, 2B, and 6B.
@@
StarCoderBase [24] is jointly released by Hugging Face and ServiceNow in 2023. It is trained on 80+ programming languages from The Stack [72]. And the renowned StarCoder is a fine-tuned version of StarCoderBase, trained on an additional 35 billion Python tokens. Considering that StarCoder is only available in a 16B version, which may not be suitable for conducting comprehensive comparison experiments. In our experiments, we use three sizes of StarCoderBase: 1B, 3B, and 7B.
@@
ChatGPT [21] is possibly the most powerful LLM. To conduct experiments with its assistance, we utilize the OpenAI API, which is powered by a diverse set of models with different capabilities, such as GPT-3.5, DALL·E and Whisper. In our experiments, we use the gpt-3.5-turbo model, which distinguishes itself as the most capable and cost-effective option within the GPT-3.5 family.
4.1.4 Experimental Settings
@@
At the training stage, we set the mini-batch size to 16, and the learning rate to 5e-5. We employ the AdamW optimizer [73] along with a linear learning rate scheduler, following the default setup recommended by Hugging Face. To ensure comprehensive training, we employ an early stopping strategy based on the best validation BLEU to control the number of training epochs for the model. The early stopping patience is set to 4. The input sequences are padded to the maximum length with special tokens, which are selected based on the vocabulary of the LLM. To fine-tune StarCoderBase-7B and CodeGen-Multi-6B on a single A800, we utilize DeepSpeed 0.12.2. We set the Zero Redundancy Optimizer (ZeRO) to ZeRO-3 and enable the offloading of optimizer computation to CPU. All models are implemented using the PyTorch 2.1.0 framework with Python 3.8. All experiments are conducted on a server equipped with one Nvidia A800 GPU with 80 GB memory, running on Ubuntu 20.04.4.
4.2 Experimental Results
4.2.1 RQ1: Effectiveness of PromptCS
Instruct Prompting (zero-shot). This scheme directly uses human-written instructions to prompt LLMs to generate summaries for given code snippets. For StarCoderBase, CodeGen-Multi, and PolyCoder, we utilize the human-written prompt provided in the work [74]. For ChatGPT, we utilize the human-written prompt provided in the work [5]. Table 1 presents the two human-written prompts mentioned above. The prompts we use have been proven effective in previous works. Considering that different prompts would threaten the fairness of RQ1, we also tried both prompt instructions (PI1 and PI2) for each LLM.
Table 1: Manual prompts used in zero-shot learning to guide LLMs for code summarization
Number LLM Prompt Instruction
[…]
CodeGen-Multi
PolyCoder
[…]
Instruct Prompting (few-shot). In this scheme, we provide a few examples that demonstrate the nature of the task to let LLMs perform few-shot learning. We follow [28, 74] and provide 10 examples in the few-shot setting. Each example is a pair of $\langle code\;snippet,summary\rangle$ randomly selected from the training set. In practice, we directly leverage the 10 examples provided by Ahmed et al. in their GitHub repository [75], since we use the same experimental dataset (i.e., the CSN corpus).
@@ 4.2.1 RQ1: Effectiveness of PromptCS
Table 2: Effectiveness of PromptCS on the CSN-Java dataset. Zero-shot: Instruction Prompting with zero-shot learning; Few-shot: Instruction Prompting with few-shot learning; $\mathcal{B}$: BLEU; $\mathcal{M}$: METEOR; $\mathcal{R}$: ROUGE-L; $\mathcal{S}$: SentenceBERT.
@@
7B 9.47 8.05 20.10 30.24 16.65 12.91 33.67 58.31 21.86 15.08 41.25 63.21
ChatGPT 9.57 14.89 20.75 54.11 14.68 15.97 34.61 59.89
  • StarCoderBase-7B and CodeGen-Multi-6B are too large for effective fine-tuning.
@@
  • ChatGPT is a non-open source model and cannot be fine-tuned. We also cannot apply PromptCS to it.
@@
2) Results: Table 2 shows the performance of our PromptCS and baselines in terms of the four evaluation metrics, i.e., BLEU, METEOR, ROUGE-L, and SentenceBERT. Observe that task-oriented fine-tuning performs best on the LLM PolyCoder in terms of scores of the four metrics, followed by PromptCS, instruction prompting (few-shot), and instruction prompting (zero-shot). PromptCS outperforms zero-shot learning and few-shot learning, indicating that continuous prompts produced by the well-trained prompt agent are better than human-written prompts. The superior performance of task-oriented fine-tuning can be attributed to its approach of fully tuning model parameters on downstream task datasets, enabling the model to adapt to task-specific patterns, vocabulary, and objectives. This enhances relevance and accuracy while minimizing interference from irrelevant data. On the LLMs CodeGen-Multi and StarCoderBase, PromptCS is overall comparable to task-oriented fine-tuning, and both are significantly better than the two instruction prompting schemes. PromptCS even outperforms task-oriented fine-tuning on some LLMs, such as CodeGen-Multi-2B (row 7), StarCoderBase-1B (row 9) and -3B (row 10). We also conduct instruction prompting schemes on a commercial LLM ChatGPT, and its performance is shown in the last row of Table 2. Observe that the performance of both instruction prompting schemes on ChatGPT is better than that on the other three LLMs in all four metrics, but is also not satisfactory compared to task-oriented fine-tuning and PromptCS. Although this comparison is unreasonable, we have reason to believe that if ChatGPT is open source, the task-oriented fine-tuning scheme and PromptCS will significantly improve its code summarization capabilities compared to instruction prompting schemes.
Table 3: LLMs’ zero-shot learning performance using PI1 and PI2.
Base Model Model Size BLEU METEOR ROUGE-L SentenceBERT
PI1 PI2 PI1 PI2 PI1 PI2 PI1 PI2
PolyCoder 160M 7.98 4.72 3.40 3.76 13.82 5.99 16.12 38.17
0.4B 7.18 6.80 2.67 5.54 11.55 12.33 16.89 38.03
2.7B 7.68 4.99 2.28 3.95 12.23 7.40 12.92 35.06
CodeGen-Multi 350M 6.13 4.17 7.46 3.51 13.76 6.0 40.26 34.13
2B 7.96 3.14 3.08 2.55 13.42 5.16 15.26 23.94
6B 8.00 3.89 4.19 3.50 14.59 7.48 19.63 26.46
StarCoderBase 1B 5.95 8.29 4.27 7.8 11.52 18.42 24.68 36.00
3B 9.64 6.95 8.64 6.42 20.70 15.4 33.93 27.75
7B 9.47 5.57 8.05 4.70 20.10 12.36 30.24 20.86
ChatGPT - 3.39 9.57 12.48 14.89 12.76 20.75 49.48 54.11
To eliminate the threat of different prompts to the fairness of experimental results, we use both PI1 and PI2 for zero-shot learning on each LLM. The results are shown in Table 3. It is observed that ChatGPT performs better when prompting with PI2. While for other LLMs, there is no unified conclusion about whether PI1 or PI2 is a more effective prompt, depending on the base LLM and the metric. This again reflects the difficulty of designing a high-quality prompt instructions. But in general, for LLMs except for ChatGPT, whether using PI1 or PI2, the performance of zero-shot learning does not exceed task-oriented fine-tuning and PromptCS. Thus, it does not threaten the rationality of our conclusion.
@@
Table 4: Comparison of task-oriented fine-tuning and PromptCS in training and inference time.
@@
Training Time Inference Time Training Time Inference Time
CodeGen-Multi 350M 20h:06m:44s 02h:16m:03s 15h:01m:39s 00h:48m:45s
2B 81h:08m:24s 09h:33m:44s 58h:43m:29s 04h:15m:24s
6B 151h:55m:57s 17h:26m:00s 88h:03m:40s 08h:34m:52s
PolyCoder 160M 12h:25m:19s 01h:09m:04s 09h:41m:45s 00h:26m:44s
0.4b 22h:18m:34s 02h:25m:45s 14h:32m:49s 00h:50m:47s
2.7b 81h:54m:24s 09h:24m:08s 27h:41m:15s 02h:12m:37s
StarCoderBase 1B 50h:06m:52s 03h:42m:50s 29h:17m:32s 01h:33m:52s
3B 82h:26m:21s 08h:53m:41s 43h:51m:44s 02h:11m:46s
7B 211h:05m:18s 20h:58m:07s 67h:21m:31s 03h:56m:10s
  • StarCoderBase-7B and CodeGen-Multi-6B are too large for effective fine-tuning. Here is the single-epoch training time for it.
In addition, we also compare the training costs between the task-oriented scheme and PromptCS in terms of training time. Table 4 lists the time costs incurred during training and inference on three LLMs. Observe that compared to the task-oriented scheme, PromptCS noticeably requires less time on all LLMs, which can be attributed to the fact that PromptCS is non-invasive to LLMs and does not update the parameters of the LLMs during training. It should be noted that when the model size of the LLM is very large (e.g., StarCoderBase-7B), the task-oriented fine-tuning will be very time-consuming and become unacceptable if the available training resources (e.g., the performance and number of GPU devices) are few. In terms of inference time, the inference time of PromptCS is much shorter than that of the corresponding task-oriented fine-tuned LLM. The reason behind this is that PromptCS is non-invasive to the parameters of LLMs, so LLMs’ weights can be loaded with half precision (fp16 instead of fp32), which speeds up the inference and reduces GPU usage at the cost of only subtle accuracy loss. Hence, except for the inference time, the memory usage of PromptCS is also much less than the memory usage of task-oriented fine-tuning. For example, when the base LLM is PolyCoder-0.4b, fine-tuning the LLM requires about 11.5GB while PromptCS only takes up 2.5GB.
@@
Table 5: Effects of prompt length and concatenation mode on PromptCS. $\mathcal{B}$: BLEU; $\mathcal{M}$: METEOR; $\mathcal{R}$: ROUGE-L; $\mathcal{S}$: SentenceBERT.
@@
4.2.2 RQ2. Influence of key configurations on PromptCS
@@
Table 6: Performance of PromptCS when building the prompt encoder on different network architectures. We bold the metric scores where the prompt encoder built on BiLSTM/Transformer performs better.
@@
4.2.3 RQ3. Influence of the network architecture used in the prompt encoder on PromptCS
@@
4.2.4 RQ4. Influence of the training data size on PromptCS
In this paper, we also analyze the impact of training data size on the effectiveness of PromptCS and the task-oriented fine-tuning. To disclose this impact, we commence with a small-scale training set and systematically augment the number of training samples. The smaller training sets are randomly sampled from the complete training data. The considered training set sizes encompass 100, 1000, 10000, and 164923 (complete training data). In this experiment, we also uniformly utilize StarCoderBase-1B as the base LLM and set the prompt length and the concatenation mode to 100 and the back-end mode, respectively.
Figure 4: Training costs for different prompt lengths
Figure 5: Influence of training set size on task-oriented fine-tuning
Figure 6: Influence of training set size on PromptCS
Figure 6 and Figure 6 showcase the experimental results, where the x-axis denotes the training set size, and the y-axes denote BLEU, METEOR, ROUGE-L, and SentenceBERT. Observe that the LLM performs significantly better when adopting the task-oriented fine-tuning on the complete dataset. Contrastly, with an increase in the size of the training set, the scores of PromptCS on the four metrics will increase, but the increase is not obvious. And PromptCS trained with 100 samples performs comparably to PromptCS trained with 164,923 samples. This underscores PromptCS’s superior adaptability and generalization capabilities on a small-scale dataset. It holds profound practical significance for reducing training costs, improving training efficiency, and obtaining satisfactory performance in environments with limited data.
@@
Table 7: Effectiveness of PromptCS in different programming languages. $\mathcal{B}$: BLEU; $\mathcal{M}$: METEOR; $\mathcal{R}$:ROUGE-L; $\mathcal{S}$: SentenceBERT. We bold the metric scores where Task-oriented Fine-tuning/PromptCS performs better.
@@
Task-oriented Fine-tuning PromptCS Task-oriented Fine-tuning PromptCS
@@
4.2.5 RQ5. Effectiveness in other programming languages